Which equation is equivalent to
4(2-5x)=6-3(1-3x)

Answer by stanbon(75887)

Which equation is equivalent to
4(2-5x)=6-3(1-3x)
8-20x = 6-3+9x
8-20x = 3+9x
29x = 5
x = 5/29
=============
Cheers,
Stan H.